# thanks: Part of the Instituto Avanzado de Cosmología (IAC) Collaboration

### v1

@@ IV.3 Centaurus A
In this subsection, we analyze the galaxy Centaurus A (also known as NGC 5128). The galaxy is an active galactic nucleus, with a supermassive black hole at its center ejecting relativistic jets emitting in X-ray and radio wavelengths. Again, we exclude this innermost region and the region where the bulge is dominant, and fit a exponential disc only.
@@
We use the rotation curve obtained in (Struve et al., 2010), with ATCA 21-cm line observations of the neutral hydrogen (HI). The HI kinematics shows that the galaxy has a regularly rotating, warped main disc (for $r<6$ kpc). No bar or bulge is necessary to reproduce the HI kinematics. The total extent of the HI disc is $\sim 7.5$ kpc in radius, and its total mass is $4.9\times 10^{8}M_{\odot}$.
The best-fit parameters, obtained from the non-linear least-squares method, are listed in Table 3. Figure 6 shows the resulting best-fit rotation curve with disc and ground plus first excited mSFDM states, and the behavior of the last two densities. Notice that the best-fit rotation curve agrees well to the observations.
Disc $m_{d}=(1.5\pm 0.6)\times 10^{8}M_{\odot}$
$h_{d}=(0.144\pm 0.056)$ kpc
mSFDM $l=(1.79\pm 0.054)\text{kpc}^{-1}$
$\rho_{0}=(596.8\pm 36.2)\times 10^{-2}M_{\odot}\text{pc}^{-3}$
$\rho_{1}=(200.0\pm 18.4)\times 10^{-2}M_{\odot}\text{pc}^{-3}$
Table 3: Best-fit results for Centaurus A, from the least-squares method.
Figure 6: (Top) Rotation curve of the Centaurus A galaxy using the ground and first excited states of the mSFDM model. (Bottom) The evolution of the ground state and first excited state densities using parameter values that fit well to the Centaurus A rotation curves.
IV.4 MW-like Galaxies with Satellites
@@
In Nashimoto et al. (2022), the authors analyzed observations from Subaru/Hyper Suprime-Cam of satellite galaxies around 9 Milky Way-like galaxies located outside of the Local Group. After screening to eliminate false positives, they identified 51 dwarf satellite galaxies within the virial radius of the host galaxies.
@@
The study found that the average luminosity function of the satellite galaxies is consistent with that of the Milky Way satellites, though the luminosity function of each host galaxy varies significantly. They also found that more massive host galaxies tend to have a larger number of satellites. The physical properties of the satellites, such as the size-luminosity relation, were found to be consistent with those of the Milky Way satellites.
@@
The study concluded that this difference in spatial distribution might represent a peculiarity of the Milky Way satellites. Further research is required to understand the origin of this trend, considering that the observation was focused on relatively massive satellites with $M_{V}<-10$.
@@
Despite not observing a clear alignment of satellites in those galaxies, we used their reported observations to test the SFDM model with two quantum states and see if there is any indication of a distribution around the atom-like structure of the SF model. While the absence of a pronounced alignment of satellite galaxies might suggest that the model’s atom-like structure is not immediately evident, it is relevant to explore this possible implication of the model.
@@
Thus, we fit the rotation curves of 6 of the 9 galaxies we found, to obtain the best parameters of the model, with mSFDM only. Remarkably, all the galaxies could be successfully reproduced with the ground and the first excited states only, corresponding to 3 free parameters. In this case, we implemented a Markov Chain Monte Carlo (MCMC) code with the emcee library in Python. We report the errors at $1\sigma$. The results of these fits are presented in Table 4. Figures 7-9 show the best-fit rotation curves for each galaxy, and the figures reproduced from Nashimoto et al. (2022) for NGC 3338, 3437, 5866 and 7332, showing the central galaxy and the confirmed and tentative satellites within the virial radius of each galaxy. We superposed the two-lobes structure from our best fits for the host galaxies, with radius $R_{\text{max}}^{1}$ (yellow solid circles).
@@
Although the alignment in all the galaxies in (Nashimoto et al., 2022) is not clear, in some cases it would be possible to explain that dwarf galaxies are found within the gravitational potential traced by the two-lobed structure of the first excited state, as in NGC 3338 (the galaxies near the center), two dwarf galaxies in NGC 3437 near the center, around 7 satellite galaxies in NGC 5866, and one or two galaxies around NGC 7332. The corresponding figures for NGC 2950 and NGC 3245 can be found in Tanaka et al. (2018); in these cases, for NGC 2950 is possible to found 3 galaxies inside the gravitational potential of the lobes, and in NGC 3245 about 6 galaxies.
Host Galaxy $\rho_{0}$ $\rho_{1}$ $l$ $R_{0}^{\text{max}}$ $R_{1}^{\text{max}}$ $R_{\text{vir}}$
($10^{-2}M_{\odot}\text{pc}^{-3}$) ($10^{-2}M_{\odot}\text{pc}^{-3}$) ($\text{kpc}^{-1}$) (kpc) (kpc) (kpc)
MW $10.04\pm 4.27$ $25.79\pm 12.63$ $0.511\pm 0.072$ 120.0 190.0 385.0
M 31 $28.60\pm 8.53$ $12.12\pm 9.09$ $0.495\pm 0.029$ 206.0 133.0 270.0
Cen A $596.8\pm 36.2$ $200.0\pm 18.4$ $1.79\pm 0.054$ 269.5 152.5 409.0
NGC 2950 $1203\pm 107$ $358\pm 30$ $4.48\pm 0.40$ 154.0 78.5 176.0
NGC 3245 $145.9\pm 56.2$ $432.8\pm 293.3$ $2.81\pm 0.80$ 84.0 145.0 226.7
NGC 3338 $0.798\pm 0.045$ $1.134\pm 0.413$ $0.133\pm 0.014$ 130.0 140.5 266.4
NGC 3437 $182.5\pm 7.91$ $185.9\pm 22.59$ $2.146\pm 0.081$ 122.0 124.5 243.0
NGC 5866 $52.81\pm 4.48$ $306.3\pm 14.4$ $1.562\pm 0.030$ 91.5 217.0 301.2
NGC 7332 $24.42\pm 2.91$ $30.72\pm 17.15$ $0.905\pm 0.142$ 109.0 122.0 210.0
Table 4: Results for the mSFDM halos. The virial radius for the MW is taken from (Sofue, 2012), for M31 is from (Corbelli et al., 2010), and for Cen A is taken from (Müller et al., 2019). The virial radius for the last 6 galaxies is taken from Nashimoto et al. (2022).
@@
Figure 7: Rotation curves of galaxies NGC 2950 and NGC 3245 using the ground state and the first excited state of the SFDM model. Note that the fit of the observed rotation curves agrees excellently with the model.
@@
Figure 8: (Left) Rotation curves of galaxies NGC 3388 and NGC 3437 using the ground state and first excited state of the mSFDM model. The fit of the observed rotation curves agrees excellently with the model. (Right) The same galaxies showing their satellite galaxies reproduced from the reference Nashimoto et al. (2022), with permission of the authors. The red and blue circles represent the secure and possible dwarf galaxies, respectively. The green shaded circles and boxes show the masked regions. The white big circle is the virial radius of the host galaxy. In each galaxy, we superposed the yellow solid circles, which show the two-lobes excited state of the mSFDM.
@@
Figure 9: (Left) Rotation curves of galaxies NGC 5866 and NGC 7332 using the ground state and first excited state of the SFDM model. The fit of the observed rotation curves agrees excellently with the model. (Right) The same galaxies showing their satellite galaxies reproduced from the reference Nashimoto et al. (2022), with permission of the authors. The red and blue circles represent the secure and possible dwarf galaxies, respectively. The green shaded circles and boxes show the masked regions. The white big circle is the virial radius of the host galaxy. In each galaxy, we superposed the yellow solid circles, which show the two-lobes excited state of the mSFDM.
@@
The whole analysis in this Section confirms that the SFDM model with two states is capable of accurately describe the rotation curves of the galaxies under study. The fact that all galaxies were well-fitted with the same number of parameters further supports the robustness and universality of the model in capturing the essential features of the observed rotation curves.
@@
This is a significant advantage over the finite temperature model with real SF presented in Robles and Matos (2013a), which required various combinations of ground plus excited states to reproduce different observations. In our model with a complex SF, using the ground state and the first excited state it possible to reproduce all the rotation curves of the studied galaxies. In other words, with the spherical structure of the ground state and the two-lobed structure of the first excited state, it is sufficient to reproduce the observations and also explain the gravitational potential responsible to align the satellite galaxies as observed in MW, M31, and CenA, and possibly the galaxies in (Nashimoto et al., 2022).
V Conclusions
@@
The SFDM model has been very successful in explaining several characteristics of DM, which other models cannot explain or need extra physics to do so. The field equations of the SFDM are the system of Schrödinger-Poisson equations. In this work, we consider the quantum character of this model to explain the non-homogeneous distribution of satellite galaxies around their host galaxies. This phenomenon has been observed in the MW, Andromeda, and Centaurus A galaxies, the first where the behavior of the satellite galaxies can be well seen. We solve the Schrödinger equation using separation of variables, where the angular dependence is given by the spherical harmonic functions, which contain terms axial symmetrical. We use these terms to model the inhomogeneous characteristic of the DM and fit the free parameters of the solutions using the rotation curves of the host galaxies. With this, we can explain the anisotropic distribution of satellite galaxies around their host galaxy, with the two-lobes shape of the first excited state of the multistate SFDM, with complex SF within the finite temperature scenario.
@@
We also find that the scale $l$ of the host galaxy depends on the energy of the vibrations $\omega$ and the effective mass $m_{\Phi}$. The SFDM collapses after spinning and this collapse causes the temperature of the SFDM to rise again. The final temperature therefore determines the effective mass of the SF, and this mass determines the scale of the galaxy. This is why each galaxy has different $l$ scales. Using the values found through the rotation curves, we can obtain the size of the excited states that can explain the VPOS in these galaxies. The fact that the scale $l$ depends on the galaxy we are studying explains why some observations fit the SFDM mass to one value and different observations of galaxies give a different value, since the scale is not determined by the SF mass, but by the scale $l$ that depends on the final collapse temperature of each galaxy. This is the first natural explanation for the anomalous behavior of satellite galaxies, with the surprising result that the first three galaxies where satellite galaxies can be seen with good resolution, have the same behavior. If in the future we can see more galaxies where their satellite galaxies maintain the same behavior, this will be strong support for the SFDM paradigm.

### v3

@@ 5.3 Centaurus A
Centaurus A (NGC 5128) is an active galactic nucleus, with a supermassive black hole at its center ejecting relativistic jets emitting in X-ray and radio wavelengths. It is a peculiar galaxy since it is in collision with another spiral galaxy. For the analysis, we exclude the innermost region and the region where the bulge is dominant, thus we fit an exponential disc with the mSFDM halo.
@@
We use the rotation curve obtained in [73], with ATCA 21-cm line observations of the neutral hydrogen (HI). The HI kinematics shows that the galaxy has a regularly rotating, warped main disc (for $r<6$ kpc). No bar or bulge is necessary to reproduce the HI kinematics. The total extent of the HI disc is $\sim 7.5$ kpc in radius, and its total mass is $4.9\times 10^{8}M_{\odot}$.
5.4 MW-like Galaxies with Satellites
@@
In [65], the authors analyzed observations from Subaru/Hyper Suprime-Cam of satellite galaxies around 9 Milky Way-like galaxies located outside of the Local Group. After screening to eliminate false positives, they identified 51 dwarf satellite galaxies within the virial radius of the host galaxies.
@@
The study found that the average luminosity function of the satellite galaxies is consistent with that of the Milky Way satellites, though the luminosity function of each host galaxy varies significantly. They also found that more massive host galaxies tend to have a larger number of satellites. The physical properties of the satellites, such as the size-luminosity relation, were consistent with those of the Milky Way satellites.
@@
The study concluded that this difference in spatial distribution might represent a peculiarity of the Milky Way satellites and further research is required to understand the origin of this trend, considering that the observation was focused on relatively massive satellites with $M_{V}<-10$.
@@
Despite not observing a clear alignment of satellites in those galaxies, we used their reported observations to test the SFDM model with two quantum states and see if there is any indication of a distribution around the atom-like structure of the SF model. While the absence of a pronounced alignment of satellite galaxies might suggest that the model’s atom-like structure is not immediately evident, exploring this possible implication of the model is relevant.
Thus, we fit the rotation curves of 6 of the 9 galaxies we found to obtain the best parameters of the model, with a mSFDM halo.
The rotation curve and the data for NGC 2950 are taken from [74]. The line-of-sight velocity data for NGC 3437 are taken from [77], from which we made the interpolation between the approaching and the receding parts of the galaxy to obtain the rotation curve and errors we used in the present work. The position-velocity map for NGC 3338 is taken from [76], from which we obtained the rotation curve and errors used in this work. The rotation velocities for NGC 3245 and NGC 7332 are taken from [75] and [79], respectively. For NGC 5866 the rotation velocities are taken from [78], from which we interpolated between the approaching and the receding parts of the galaxy.
6 Results
Remarkably, all the galaxies could be successfully reproduced with the ground and the first excited states of the mSFDM model studied in this work. The best-fit parameters, obtained from the MCMC method implemented in Python (see Section 4), are listed in table 2. The errors are reported to $1\sigma$ of confidence level. We also show the posterior distributions for the corresponding parameters (figures 4-7).
6.1 Milky Way
The best-fit parameters, obtained for the exponential bulge and disc, and the ground and first excited states of the mSFDM halo, are listed in table 2.
In figure 4, we show the best-fit MW rotation curve up to 64.5 kpc and the densities of the ground and first excited states of the mSFDM configuration. We cut the DM contribution as defined in (2.33). The posterior distributions for the 7 free parameters are also shown in this figure.
Figure 4: (Top) Best fit results for the Milky Way with bulge, disc, and mSFDM components, in normal and logarithmic scale, the last one showing the central region of the Galaxy. (Bottom) The behavior of the densities, using the same best-fit parameters. Posterior distributions for the parameters $\rho_{0}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $\rho_{1}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $l(\text{kpc}^{-1})$, $m_{b}(10^{10}M_{\odot})$, $h_{b}$(kpc), $m_{d}(10^{10}M_{\odot})$, and $h_{d}$(kpc). The contour lines correspond to the $1\sigma$, $2\sigma$, and $3\sigma$ confidence regions.
In table 3, we compare the maximum radius of the mSFDM configuration, with the virial radius reported in [82]. Both radii, $R_{0}^{\text{max}}$ and $R_{1}^{\text{max}}$, are around 0.3 and 0.5 times the virial radius, respectively. With this result, the two-lobe structure, originating in the center of the Milky Way and orthogonal to its plane, would enclose around 53 of 61 satellites listed in [83].
6.2 Andromeda
The best-fit parameters for Andromeda for an exponential disc and mSFDM halo are listed in table 2. In figure 5, we show the best-fit M31 rotation curve up to 36.5 kpc and the densities of the ground and first excited states. We also show the posterior distributions for the 5 free parameters.
Figure 5: (Top) Best-fit rotation curve of M31 with disc and mSFDM components. Densities of the mSFDM ground and first excited state for the same best-fit parameters. (Bottom) Posterior distributions for the parameters $\rho_{0}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $\rho_{1}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $l(\text{kpc}^{-1})$, $m_{d}(10^{10}M_{\odot})$, and $h_{d}$(kpc). The contour lines correspond to the $1\sigma$, $2\sigma$, and $3\sigma$ confidence regions.
In table 3, we show the maximum radii of both mSFDM states, the ground and the first excited states, and the virial radius reported in [72]. Such radii are around 0.75 and 0.5 times the virial radius, respectively. With this result, the two-lobe structure, originating in the center of Andromeda and orthogonal to its plane, would enclose 10 of 18 satellites listed in [84].
6.3 Centaurus A
The best-fit parameters, obtained from the MCMC method for an exponential disc and mSDFM halo, are listed in table 2. Figure 6 shows the resulting best-fit rotation curve and the behavior of the SF densities. Notice that the best-fit rotation curve agrees well with the observations. We also show the resulting posterior distributions for the 5 free parameters.
Figure 6: (Top) Best-fit rotation curve of Centaurus A with disc and mSFDM components. Densities of the mSFDM ground and first excited state for the same best-fit parameters. (Bottom) Posterior distributions for the parameters $\rho_{0}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $\rho_{1}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $l(\text{kpc}^{-1})$, $m_{d}(10^{10}M_{\odot})$, and $h_{d}$(kpc). The contour lines correspond to the $1\sigma$, $2\sigma$, and $3\sigma$ confidence regions.
In table 3, we show the maximum radii of both mSFDM states, the ground and the first excited states, and the virial radius reported in [85]. Such radii are around 0.7 and 0.4 times the virial radius, respectively. With this result, the two-lobe structure, originating in the center of Centaurus A and orthogonal to its plane, would enclose 15 of 31 satellites (5 with velocity measurements and 10 without velocities), and mainly the whole structure of 1239 planetary nebulae reported in [10] (see the central squared region of figure 1 in that work).
6.4 MW-like Galaxies with Satellites
@@
For these galaxies, we used a mSFDM halo and only for NGC 3437 we included an exponential disc to better fit the data. The results of these fits are presented in table 2. Figures 8-10 show the best-fit rotation curves for each galaxy, and the figures reproduced from [65] for NGC 3338, 3437, 5866 and 7332, show the central galaxy and the confirmed and tentative satellites within the virial radius of each galaxy. We superposed with yellow solid circles the resulting two-lobe structure (the excited state) from our best fits for the host galaxies, with radius $R_{\text{max}}^{1}$ (see table 3).
@@
Although the alignment in all the galaxies in [65] is not clear, in some cases it would be possible to explain that dwarf galaxies are found within the gravitational potential traced by the two-lobe structure of the first excited state, as in NGC 3338 (the galaxies near the center), two dwarf galaxies in NGC 3437 near the center, about 7 satellite galaxies in NGC 5866, and one or two galaxies around NGC 7332. The corresponding figures for NGC 2950 and NGC 3245 can be found in [86]; in these cases, for NGC 2950 it is possible to find 3 galaxies inside the gravitational potential of the lobes, and in NGC 3245 about 6 galaxies.
Figure 7: Posterior distributions for the parameters $\rho_{0}(10^{-2}M_{\odot}/\text{pc}^{-3})$, $\rho_{1}(10^{-2}M_{\odot}/\text{pc}^{-3})$, and $l(\text{kpc}^{-1})$, for the galaxies with mSFDM-only, NGC 2950, 3245, 3338, 5866 and 7332. In the case of NGC 3437, the contours include $m_{d}(10^{10}M_{\odot})$ and $h_{d}$(kpc), the parameters of the exponential disc. The contour lines correspond to the $1\sigma$, $2\sigma$, and $3\sigma$ confidence regions.
mSFDM halo results
Host Galaxy $\rho_{0}$ $\rho_{1}$ $l$ $M_{\text{halo}}$ $-2\ln\mathcal{L}$
($10^{-2}M_{\odot}\text{pc}^{-3}$) ($10^{-2}M_{\odot}\text{pc}^{-3}$) ($\text{kpc}^{-1}$) $(10^{12}M_{\odot})$
MW $3.88^{+1.16}_{-1.08}$ $3.19^{+1.79}_{-1.69}$ $0.23^{+0.03}_{-0.03}$ 1.38 2.7
M 31 $3.46^{+0.38}_{-0.40}$ $0.766^{+0.477}_{-0.323}$ $0.169^{+0.007}_{-0.006}$ 1.8 7.22
Cen A $750.77^{+133.42}_{-115.18}$ $356.31^{+193.57}_{-153.68}$ $2.09^{+0.20}_{-0.20}$ 3.7 3.98
NGC 2950 $1204.8^{+108.74}_{-101.07}$ $395.62^{+240.39}_{-171.03}$ $4.52^{+0.32}_{-0.30}$ 0.674 4.78
NGC 3245 $155.63^{+44.5}_{-38.0}$ $383.03^{+179.89}_{-163.9}$ $2.70^{+0.50}_{-0.56}$ 0.598 0.16
NGC 3338 $0.79^{+0.03}_{-0.04}$ $1.13^{+0.41}_{-0.41}$ $0.13^{+0.01}_{-0.01}$ 1.06 5.11
NGC 3437 $28.26^{+5.56}_{-5.42}$ $16.78^{+5.67}_{-5.93}$ $0.70^{+0.06}_{-0.07}$ 0.789 38.33
NGC 5866 $52.87^{+4.54}_{-4.58}$ $306.2^{+14.62}_{-14.72}$ $1.56^{+0.02}_{-0.02}$ 1.84 39.02
NGC 7332 $24.69^{+3.05}_{-2.78}$ $32.05^{+11.05}_{-10.96}$ $0.91^{+0.08}_{-0.09}$ 0.502 13.48
Baryonic components results
Disc Bulge
Host Galaxy $m_{d}$ $h_{d}$ $m_{b}$ $h_{b}$
($10^{10}M_{\odot}$) (kpc) ($10^{10}M_{\odot}$) (pc)
MW $4.53^{+0.32}_{-0.34}$ $2.58^{+0.23}_{-0.17}$ $1.08^{+0.09}_{-0.06}$ $0.17^{+0.02}_{-0.01}$
M31 $6.13^{+1.02}_{-0.67}$ $4.59^{+0.81}_{-0.91}$
Cen A $0.012^{+0.001}_{-0.001}$ $0.10^{+0.032}_{-0.035}$
NGC 3437 $0.34^{+0.07}_{-0.06}$ $0.48^{+0.05}_{-0.04}$
Table 2: Best-fit results from the MCMC method. (Top) Results for the mSFDM halos reported with $1\sigma$ errors. The halo mass, $M_{\text{halo}}$, is calculated from the sum of the ground and excited states, up to $R^{\text{max}}$, reported in table 3. We report the minimum values $-2\ln\mathcal{L}$ from the MCMC method. (Bottom) Results of the baryonic components reported with $1\sigma$ errors. To fit the MW rotation curve, we use an exponential bulge and disc. For M31, Cen A, and NGC 3437, we use an exponential disc.
Radii for the mSFDM configurations
Host Galaxy $R_{0}^{\text{max}}$ $R_{1}^{\text{max}}$ $R_{\text{vir}}$ $R_{1}^{\text{max}}/R_{\text{vir}}$
(kpc) (kpc) (kpc)
MW 170.0 159.0 385.0 0.41
M 31 213.0 111.0 270.0 0.41
Cen A 257.5 180.0 409.0 0.44
NGC 2950 153.7 86.5 176.0 0.49
NGC 3245 92.2 143.9 226.7 0.63
NGC 3338 130.0 166.0 266.4 0.62
NGC 3437 149.7 116.0 243.0 0.47
NGC 5866 91.0 217.5 301.2 0.72
NGC 7332 108.0 124.0 210.0 0.59
Table 3: Radii for the mSFDM halos. The maximum radii of both mSFDM states are obtained such that $\rho_{i}(R_{i}^{\text{max}})=200\rho_{c}$ (cf. equation (2.33)), for the densities given by equations (2.31) and (2.32). The virial radius for the MW is taken from [82], for M31 is from [72], and for Cen A is taken from [85]. The virial radius for the last 6 galaxies is taken from [65]. We show also the ratio between $R_{1}^{\text{max}}$, the radius of the two-lobe structure, and $R_{\text{vir}}$.
@@
Figure 8: Rotation curves of galaxies NGC 2950 and NGC 3245 using the ground state and the first excited state of the SFDM model. Note that the fit of the observed rotation curves agrees well with the model.
@@
Figure 9: (Left) Rotation curves of galaxies NGC 3388 and NGC 3437 using the ground and first excited states of the mSFDM model. The fit of the observed rotation curves agrees well with the model. (Right) The same galaxies showing their satellite galaxies reproduced from [65], with permission of the authors. The red and blue circles in the original figures represent the secure and possible dwarf galaxies, respectively. The green shaded circles and boxes show the masked regions. The big white circle, also in the original figures, is the virial radius of the host galaxy. In each galaxy, we superposed the yellow solid circles, which show the two-lobe excited state of the mSFDM.
@@
Figure 10: (Left) Rotation curves of galaxies NGC 3388 and NGC 3437 using the ground and first excited states of the mSFDM model. The fit of the observed rotation curves agrees well with the model. (Right) The same galaxies showing their satellite galaxies reproduced from [65], with permission of the authors. The red and blue circles in the original figures represent the secure and possible dwarf galaxies, respectively. The green shaded circles and boxes show the masked regions. The big white circle, also in the original figures, is the virial radius of the host galaxy; for NGC 5866 is outside the observational region. In each galaxy, we superposed the yellow solid circles, which show the two-lobe excited state of the mSFDM.
@@
The whole analysis in this Section confirms that the mSFDM model with two states is capable of accurately describing the rotation curves of the galaxies under study. The fact that all galaxies were well-fitted with the same number of parameters further supports the robustness and universality of the model in capturing the essential features of the observed rotation curves.
@@
This is a significant advantage over the finite temperature model with real SF presented in [43], which required various combinations of ground plus excited states to reproduce different observations [43, 56, 23]. In the present model, with a complex SF and by using the spherical ground state and the two-lobe first excited state only, it is possible to reproduce all the rotation curves of the studied galaxies. In other words, with the spherical structure of the ground state and the two-lobe structure of the first excited state, it is sufficient to reproduce the observations and also explain the gravitational potential responsible for aligning the satellite galaxies as observed in MW, M31, and CenA, and possibly the galaxies in [65].
7 Conclusions
@@
The SFDM model has been very successful in explaining several characteristics of DM, which other models cannot explain or need extra physics. The field equations of the SFDM are the system of Schrödinger-Poisson equations. In this work, we consider the quantum character of this model to explain the non-homogeneous distribution of satellite galaxies around their host galaxies. This phenomenon has been observed in the Milky Way, Andromeda, and Centaurus A, where the behavior of the satellite galaxies has been well studied. We started from a complex SF in thermal equilibrium with other components of the SM, at the very early stages of the Universe. Then we discuss the conditions for an SSB to take place as the Universe cools down. Then we solve the evolution equations in an FLRW Universe, separating the SF function into its background and linear perturbations, and show the Newtonian limit to find the analytic solutions for the density. Then we solve the Schrödinger equation by using separation of variables, where the angular dependence is given by the spherical harmonic functions, which contain axial symmetric terms. We use such terms to model the inhomogeneous behavior of the DM and fit the free parameters of the solutions using the rotation curves of the Milky Way, Andromeda, Centaurus A, and 6 other MW-like galaxies as the host galaxies of dwarf satellites. With this, we can explain the anisotropic distribution of satellite galaxies around their host galaxy, with the two-lobe shape of the first excited state of the multistate SFDM, from complex SF within the finite temperature scenario.
This is the first natural explanation for the anomalous behavior of the satellite galaxies, with the surprising result that the first three galaxies where satellite galaxies can be seen with good resolution, have the same behavior. If in the future we can see more galaxies where their satellite galaxies maintain the same behavior, this will be a strong support for the multistate SFDM paradigm.
@@
We also find that the length-scale $l$ of the host galaxy depends on the energy of the vibrations $\omega$ and the effective mass $m_{\Phi}$. The SFDM collapses after spinning and this collapse causes the temperature of the SFDM to rise again. The final temperature therefore determines the effective mass of the SF, and this mass determines the scale of the galaxy. This is why each galaxy has different $l$ scales. As we have seen, this scale is determined by $l^{2}=\omega^{2}-m_{\Phi}^{2}a^{2}$. If we take the self-interaction constant $\lambda$ as in [50], the critical temperature $T_{c}$ is very high and the effective mass $m_{\Phi}=m$, is the mass of the scalar field. Therefore, the parameter $\omega$ is simply $\omega^{2}=l^{2}+m^{2}a^{2}\sim m^{2}$, since $l\sim 10^{-23}$eV. Using the values found with the rotation curves, we can obtain the size of the excited states that can explain the VPOS in these galaxies. The fact that the scale $l$ depends on the galaxy we are studying explains why some observations fit the SFDM mass to one value and different observations of galaxies give a different value, since the scale is not determined by the SF mass, but by the scale $l$ that depends on the final collapse temperature of each galaxy.